# Distinguishing radiation mechanisms and particle populations in blazar jets through long-term multi-band monitoring with RINGO3 and Fermi

### v1

@@ 1 Introduction
Furthermore, blazars can be sub-classified according to the rest-frame location of this lower-energy synchrotron ($\nu_{s}$) peak in their SEDs. This was first introduced by Padovani & Giommi (1995) and has been adapted by Abdo et al. (2010) for use on large samples of Fermi blazars. BL Lac sources are classified as low synchrotron peak (LSP; $\nu_{s}$ < 1014 Hz), intermediate synchrotron peak 1014 < $\nu_{s}$ < 1015 Hz and high synchrotron peak (HSP; $\nu_{s}$ >1015 Hz). FSRQs are all classified as LSPs, based on the location of their synchrotron peak (Abdo et al., 2010).
The origin of the higher-energy peak, located at hard X-rays to very-high-energy (VHE) $\gamma$-rays, is still debated. Leptonic modelling of this high-energy emission suggests inverse-Compton scattering as the likely origin (Maraschi et al., 1992; Bloom & Marscher, 1996; Böttcher et al., 2013). In this scenario, low-energy photons originating from within the jet (synchrotron-self Compton (SSC); Maraschi et al. (1992)) or from outside (external Compton (EC); Dermer & Schlickeiser (1993)) are upscattered by interactions with the population of synchrotron electrons within the jet, thus producing the observed high-energy emission. Conversely, assuming hadronic modelling, the acceleration of protons to VHE can lead to the high-energy emission directly through proton synchrotron emission, or via interactions between the protons producing both charged and neutral pions (Mannheim & Biermann, 1992; Aharonian, 2000; Böttcher et al., 2013). It is the decay of these charged pions that produces high-energy neutrinos, the smoking-gun signature of hadronic emission (Reimer, 2012), that have been detected coinciding with several flaring blazars (Plavin et al., 2023).
@@
A frequent optical photometric feature of blazars is their colour evolution during various levels of jet activity. Most are ‘bluer-when-brighter’ (BWB) i.e., their SED at optical frequencies flattens during periods of higher flux. This behaviour can be explained with a one-component synchrotron model with an injection of fresh electrons into the jet with a hard energy distribution. These electrons cool and the resulting increased radiation is bluer in colour. Two-component modelling suggests two underlying components to the observed flux, one stable and one variable (Fiorucci et al., 2004). The stable component consists of thermal emission from the accretion disc and broad line regions (BLR) whereas the variable component originates from non-thermal synchrotron emission. In some sources, predominantly FSRQs (Zhang et al., 2015; Negi et al., 2022), we see the opposite behaviour; that is the source appears ‘redder-when-brighter’ (RWB), or their SED at optical frequencies steepens during periods of higher flux. This is thought to be due to an increased amount of thermal emission from the disc, resulting in the composite spectrum being flatter in the optical region (the ‘blue/UV bump’; Gu et al., 2006) and subsequently steepening during periods of heightened flux. Moreover, a stable-when-brighter has been observed in some objects (Ghosh et al., 2000; Zhang et al., 2015), where the colour of the source remains constant during flux increases.
@@ 2 Observations
The RINGO2 (Steele et al., 2006) and DIPOL blazar monitoring campaign (Jermak et al., 2016), ran from 2008 to 2012 on the Liverpool Telescope. It was designed to provide optical photo-polarimetric monitoring of 15 $\gamma$-ray flaring blazars (also monitored at high energies by the Fermi Gamma-ray Space Telescope). The sample has slowly grown since 2008 with the introduction of new sources that have exhibited $\gamma$-ray flaring. With the introduction of RINGO3 (Steele et al., 2010) in 2013, photo-polarimetric monitoring of the existing sample was expanded with an additional 16 blazars. RINGO3 operated on the Liverpool Telescope until 2020; we present the photometric results of this monitoring campaign here (polarimetric results will be presented in a subsequent paper).
@@ 2.1 RINGO3 polarimeter
The Liverpool Telescope (LT) is a 2.0-m, fully autonomous, robotic telescope located on the Canary Island of La Palma (Steele et al., 2004). The LT’s intelligent dispatch scheduler allows the telescope to operate entirely autonomously, selecting observational sequences according to weather conditions, science aims, source visibility and location on sky, along with priority gradings. The LT’s autonomous operation makes it ideal for blazar monitoring as observations can be filtered in amongst other science targets regularly across periods of months, along with the possibility of intensive periods of intra-night monitoring.
@@
The three optical wavebands in RINGO3 are dictated by the dichroic mirrors. The mirrors were selected at the time of construction to maximise the amount of flux detected in gamma-ray burst follow-up and as such do not correspond with standard astronomy passbands (e.g. Johnson-Cousin or $ugriz$; Arnold et al. (2012)). In this paper the different bands will be referred to as as b* (350-640 nm), g* (650-760 nm) and r* (770-1000 nm).
@@ 2.2 Photometric calibration
[…]
The average uncertainty on each calibration star magnitude achieved was 0.07, with a standard deviation of 0.04. We note that this uncertainty in the calibration star magnitudes does not affect the correlation statistics presented in this work as it will offset all data blazar points by the same amount.
@@ 2.3 Fermi data
Fermi data were taken from the Fermi Large Area Telescope (LAT) Light Curve Repository (LCR) [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (Abdollahi et al., 2023). This database consists of flux-calibrated light curves from over 1500 variable sources (Kocevski et al., 2021) with variability indexes $>21.67$. The data can be downloaded at different binning intervals (three days, one week, and one month). We chose to use data binned over three days to ensure sufficient data when correlated against the optical.
@@ 2.4 Sample
In our program, photopolarimetric data were obtained from 49 sources between 2013 and 2020. To ensure enough data per source to measure long-term variability characteristics, a minimum of 60 observations were required for the source to be a part of our final sample. An exception was made for any object where observations were taken with a density greater than once every ten days, given that at this observation density, any correlations might be detectable. Applying these conditions resulted in 31 objects in the sample; see Table 1 for details of redshift, spectral peak (emission) classification, r*-band magnitude range and Fermi $\gamma$-ray flux range for the given observational period. This sample includes the 15 sources in the DIPOL and RINGO2 sample (Jermak et al., 2016), with additional sources selected for $\gamma$-ray activity.
@@
Name Type z r* mag. range Fermi range (ergs cm-2 s-1) Observational Period (MJD)
@@ 3 Correlation analysis
We use the Spearman rank coefficient to test the strength and significance of monotonic relationships within our data in a non-parametric way. This test provides a coefficient for significance, (p; how likely the data are to be falsely correlated) and strength (c; how strongly correlated the data are and in the positive or negative direction). Table 2 shows the correlation strength coefficient values (c) used in this analysis with $p\leq 0.05$, implying a $2\sigma$ significance level for a Gaussian distribution.
In our analysis, we account for the differing cadence of optical and $\gamma$-ray observations by assigning the temporally closest $\gamma$-ray data point to each of the optical observations. The $\gamma$-ray activity level of each source was determined by computing the median flux over the RINGO3 observation timescale. Any group of two or more points above 3 times the median absolute deviation (including lower error limits) are considered to be in a flaring state. This is indicated on the light curves in Figures 6 - 35 as a dashed purple line.
We note that no host galaxy corrections have been performed on the data. In general, blazars outshine their host galaxies by several orders of magnitude so no correction is required. We acknowledge that some sources in our sample, within the HSP BL Lac class, do have resolvable host galaxies but these are not variable. This means that for a small amount of data, a change in position on magnitude/flux axes would occur following host correction, but the correlation analysis and variability observed would remain unchanged.
@@ 3.1 Optical – gamma-ray flux
The $\gamma$-ray flux and r*-band magnitude correlations are shown in Fig. 1, with each panel showing sources of different classifications (i.e. high-, intermediate-, low- synchrotron peak BL Lac types and FSRQs) and each colour within the panel showing different sources. The data from other object classifications is shown as grey points. There is a clear distinction between the BL Lac and FSRQ sources, with some overlap between the FSRQs and BL Lac LSPs. The FSRQ sources generally have higher $\gamma$-ray fluxes however, this may just be due to the biased nature of the sample selection (sources were added to the observing campaign if they showed high levels of $\gamma$-ray activity with Fermi).
Figure 1: Fermi $\gamma$-ray flux vs optical r* magnitude for the sample. Each panel highlights the data for the different blazar subclasses separately with the upper left, upper right, lower left, and lower right highlighting data for FSRQs, BL Lac LSPs, ISP BL Lacs, and HSP BL Lacs respectively. In each panel, the data for the other subclasses is shown as faint grey circles.
Using the redshifts, $z$, from Table 1, the $\gamma$-ray fluxes and optical magnitudes can be calibrated for distance. To do this the luminosity distance, $d_{\mathrm{L}}$, was calculated for each object, in Python, using the WMAP9 cosmology module in Astropy. This module assumes a flat universe, with a Hubble constant of $H_{0}=69.32\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and the matter density parameter set at $\Omega_{\mathrm{m}}=0.2865$ (Hinshaw et al., 2013).
@@
from Hovatta et al. (2014) where $\Gamma$ is the power-law index (taken from the Fermi LCR), F is the $\gamma$-ray flux given in $\mathrm{ergs\,cm^{-2}\,s^{-1}}$ in the 1–100 GeV photon energy range, and $d_{\mathrm{L}}$ is the luminosity distance given in cm. The absolute magnitude was calculated by
@@
Figure 2: As Fig. 1, but with Fermi $\gamma$-ray flux and r* magnitude calibrated for distance ($\gamma$-ray luminosity and r* absolute magnitude respectively).
Table 5 in the appendix shows the Spearman rank p values and coefficients (c) for correlations between the RINGO3 wavebands and $\gamma$-ray flux for each source in the sample; excluding IC310, 1ES 1426+428, and MRK421 for the reasons discussed previously. Of these 28 sources, 21 showed significant correlations between $\gamma$-ray flux and each optical b*g*r* magnitude. All significant correlations were positive. Breaking these correlations down by subclassification, we find that 33 per cent of HSP BL Lac sources are significantly correlated, increasing to 66 per cent for ISP BL Lac sources. All BL Lac LSP sources showed significant correlations between the optical and $\gamma$-ray fluxes, along with 85 per cent of FSRQs.
@@ 3.2 Optical spectral index – flux
$\mathrm{F}_{\nu}\propto\nu^{-\alpha}$ (3)
where $\mathrm{F}_{\nu}$ is the flux at wavelength $\nu$, $\nu$ is the central wavelength of the RINGO3 bands, and $\alpha$ is the spectral index. A two-point spectral index was calculated by taking the gradient of log flux vs log frequency, giving $-\alpha$. This was done using the r* and b* data, and by fitting a linear least-squares regression. It is important to not include the g*-band data in this calculation to not induce false correlations in the subsequent analysis arising from correlated errors (see Appendix A for more details). The uncertainty of the spectral index at each epoch was calculated using Monte Carlo resampling where, at each epoch, 1000 pairs of randomly generated r* and b* flux values within the respective error limits were generated and the spectral index was calculated. The standard error on these 1000 values was taken as the error on the spectral index.
@@
Table 6 in the appendix shows the results of the above analysis, with the ‘Fit’ column describing which model fits the data better according to the AIC and BIC values. We note that there were no cases with conflicting AIC and BIC values. The ‘Trend’ column describes the colour relationship observed given the obtained preferred fit and sign on the Spearman rank strength coefficient, c. Negative strengths indicate bluer-when-brighter (BWB) behaviour, implying the spectral index flattens during periods of heightened flux. Conversely, positive strength coefficients indicate redder-when-brighter (RWB) behaviour implying the spectral index steepens during periods of heightened flux. In both cases, where a log fit is preferred over a linear one, the behaviour becomes more stable as brightness increases meaning the rate at which the colour changes decreases, or altogether flattens; bluer-stable-when-brighter and redder-stable-when-brighter (BSWB and RSWB, respectively). A fit line for each source is included in Fig. 3, where a dotted line indicates the linear fit was preferred or a solid line for a logarithmic fit.
@@
Figure 3: Spectral index vs optical g*-band flux for each object in our sample. A best-fit line is fitted to each set of data after having linear and logarithmic fits qualitatively assessed using AIC and BIC coefficients. A preferential linear fit is shown by a dotted fit line, while a logarithmic fit is shown by a solid line. The Spearman rank correlation coefficients associated with each correlation (linearised in the case of a logarithmic fit) are given in Table 6.

### v2

@@ 1 Introduction
Furthermore, blazars can be sub-classified according to the rest-frame location of this lower-energy synchrotron ($\nu_{\mathrm{s}}$) peak in their SEDs. This was first introduced by Padovani & Giommi (1995) and has been adapted by Abdo et al. (2010) for use on large samples of Fermi blazars. BL Lac sources are classified as low synchrotron peak (LSP; $\nu_{\mathrm{s}}$ < 1014 Hz), intermediate synchrotron peak (ISP; 1014 < $\nu_{\mathrm{s}}$ < 1015 Hz) and high synchrotron peak (HSP; $\nu_{\mathrm{s}}$ >1015 Hz). FSRQs are all classified as LSPs, based on the location of their synchrotron peak (Abdo et al., 2010).
The origin of the higher-energy peak, located at hard X-rays to very-high-energy (VHE) $\gamma$-rays, is still debated. Leptonic modelling of this high-energy emission suggests inverse-Compton scattering as the likely origin (Maraschi et al., 1992; Bloom & Marscher, 1996; Böttcher et al., 2013). In this scenario, low-energy photons originating from within the jet (synchrotron-self Compton (SSC); Maraschi et al., 1992) or from outside (external Compton (EC); Dermer & Schlickeiser, 1993) are upscattered by interactions with the population of synchrotron electrons within the jet, thus producing the observed high-energy emission. Conversely, assuming hadronic modelling, the acceleration of protons to VHE can lead to the high-energy emission directly through proton synchrotron emission, or via interactions between the protons producing both charged and neutral pions (Mannheim & Biermann, 1992; Aharonian, 2000; Böttcher et al., 2013). It is the decay of these charged pions that produces high-energy neutrinos, the smoking-gun signature of hadronic emission (Reimer, 2012), that have been detected coinciding with several flaring blazars (Plavin et al., 2023).
@@
A frequent optical photometric feature of blazars is their colour evolution during various levels of jet activity. Most are ‘bluer-when-brighter’ (BWB) i.e., their SED at optical frequencies flattens during periods of higher flux. This behaviour can be explained with a one-component synchrotron model with an injection of fresh electrons into the jet with a hard energy distribution. These electrons cool and the resulting increased radiation is bluer in colour. Two-component modelling suggests two underlying components to the observed flux, one stable and one variable (Fiorucci et al., 2004). The stable component consists of thermal emission from the accretion disc and broad line regions (BLR) whereas the variable component originates from non-thermal synchrotron emission. In some sources, predominantly FSRQs (Zhang et al., 2015; Negi et al., 2022), we see the opposite behaviour, that is: the source appears ‘redder-when-brighter’ (RWB), or its SED at optical frequencies steepens during periods of higher flux. This is thought to be due to an increased amount of thermal emission from the disc, resulting in the composite spectrum being flatter in the optical region (the ‘blue/UV bump’; Gu et al., 2006) and subsequently steepening during periods of heightened flux. Moreover, a stable-when-brighter trend has been observed in some objects (Ghosh et al., 2000; Zhang et al., 2015), where the colour of the source remains constant during flux increases.
@@ 2 Observations
The RINGO2 (Steele et al., 2006) and DIPOL blazar monitoring campaign (Jermak et al., 2016), ran from 2008 to 2012 on the Liverpool Telescope. It was designed to provide optical photo-polarimetric monitoring of 15 $\gamma$-ray flaring blazars (also monitored at high energies by the Fermi Gamma-ray Space Telescope). The sample has slowly grown since 2008 with the introduction of new sources that have exhibited $\gamma$-ray flaring. With the commissioning of RINGO3 (Arnold et al., 2012) in 2013, photo-polarimetric monitoring of the existing sample was expanded with an additional 16 blazars. RINGO3 operated on the Liverpool Telescope until 2020; we present the photometric results of this monitoring campaign here (polarimetric results will be presented in a subsequent paper).
@@ 2.1 RINGO3 polarimeter
The Liverpool Telescope (LT) is a 2.0-m, fully autonomous, robotic telescope located on the Canary Island of La Palma (Steele et al., 2004). The LT’s intelligent dispatch scheduler allows the telescope to operate entirely autonomously, selecting observational sequences according to weather conditions, science aims, source visibility and location on sky, along with priority gradings. The LT’s autonomous operation makes it ideal for blazar monitoring. Regular blazar observations can be scheduled amongst other science objects across periods of months, as well as the possibility of more intensive periods of intra-night monitoring on occasions when weather and instrument availability allow.
@@
The three optical wavebands in RINGO3 are dictated by the dichroic mirrors. The mirrors were selected at the time of construction to maximise the amount of flux detected in gamma-ray burst follow-up and as such do not correspond with standard astronomy passbands (e.g. Johnson-Cousin or $ugriz$; Arnold et al. (2012)). In this paper the different bands will be referred to as b* (350-640 nm), g* (650-760 nm) and r* (770-1000 nm).
@@ 2.2 Photometric calibration
To achieve this we used observations of unreddened A0 stars as they, by definition, have the same magnitude in all photometric passbands (i.e. zero colour) for a Vega-referenced magnitude system. A sample of bright, unreddended A0 stars with high-quality optical photometry in the Johnson-Cousins system was therefore observed with RINGO3 on a small number of photometric nights (non-coincident with the blazar observations). To account for the variable throughput of the camera/optical system over time due to dust accumulation and similar effects, approximately nightly observations of polarimetric standards (BD+64 106, BD+25 727, and HILT 960) were used to calibrate the rate of degradation. This rate was measured as $4.69\pm 0.15\times 10^{-4}$ per cent per day, irrespective of the camera, relative to the initial counts measured from the object. The counts from the A0 stars were adjusted using this degradation rate to produce zeropoints calibrated to a common date. In the same way, the counts of the calibration star in the science frames were also adjusted to the same common date. In combination with the zeropoints this allowed the calculation of the average magnitude for each calibration star per waveband in the natural RINGO3 system.
The average uncertainty on each calibration star magnitude achieved was 0.07. We note that this uncertainty in the calibration star magnitudes does not affect the correlation statistics presented in this work as it will offset all data blazar points by the same amount.
Since the electron multiplying CCD gives an effective read noise of $<1e^{-}$, we can stack images without penalty and remain photon-limited. However, electron multiplication noise reduces the final signal-to-noise ratio by a factor of $\sqrt{2}$. This means that the final photometric uncertainty is increased by this factor compared with photometry using a single conventional CCD image (Robbins & Hadwen, 2003).
@@ 2.3 Fermi data
Fermi data were taken from the Fermi Large Area Telescope (LAT) Light Curve Repository (LCR) [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (Abdollahi et al., 2023). This database consists of flux-calibrated light curves from over 1500 variable sources (Kocevski et al., 2021) with variability indexes $>21.67$. The fractional variability is described by Abdollahi et al. (2023) as a proxy for the average fractional variability exhibited by an object over a one-year timescale. The threshold of 21.67 indicates that the $\gamma$-ray flux of the object has a less than one per cent chance of being steady. The data can be downloaded at different binning intervals (three days, one week, and one month). We chose to use data binned over three days to ensure sufficient data when correlated against the optical.
@@ 2.4 Sample
In our program, photopolarimetric data were obtained from 49 sources between 2013 and 2020. To ensure enough data per source to measure long-term variability characteristics, a minimum of 60 observations were required for the source to be a part of our final sample. An exception was made for any object where observations were taken with a density greater than once every ten days, given that at this observation density, any correlations might be detectable. Application of these conditions resulted in 31 objects in the sample; see Table 1 for details on redshift, spectral peak (emission) classification, r*-band magnitude range and Fermi $\gamma$-ray flux range for the given observational period. This sample includes the 15 sources in the DIPOL and RINGO2 sample (Jermak et al., 2016), with additional sources selected for $\gamma$-ray activity.
@@
Name Type $z$ r* mag. range Fermi range (erg cm-2 s-1) Observational Period (MJD)
@@ 3 Correlation analysis
We use the Spearman rank coefficient to test the strength and significance of monotonic relationships within our data in a non-parametric way. This test provides a coefficient for significance, ($p$; how likely the data are to be correlated by chance) and strength ($c$; strength of the positive or negative correlation). Table 2 shows the correlation strength coefficient values ($c$) used in this analysis with $p\leq 0.05$, implying a $2\sigma$ significance level for a Gaussian distribution.
In our analysis, we account for the differing cadence of optical and $\gamma$-ray observations by assigning the temporally closest $\gamma$-ray data point to each of the optical observations. The median difference in MJD between optical and $\gamma$-ray observations across all sources in the sample was 1.12 days. 96 per cent of observations had a corresponding Fermi $\gamma$-ray observation within 12 days (the average cadence between optical observations across all sources). The $\gamma$-ray activity level of each source was determined by computing the median flux over the RINGO3 observation timescale. Any group of two or more points above 3 times the median absolute deviation (including lower error limits) are considered to be in a flaring state. This is indicated on the light curves in Figures 6 - 35 as a dashed purple line.
We note that no host galaxy corrections have been performed on the data. In general, blazars outshine their host galaxies by several orders of magnitude so no correction is required. We acknowledge that some sources in our sample, within the HSP BL Lac class, do have resolvable host galaxies but these are not variable. This means that for a small amount of data, a change in position on magnitude/flux axes would occur following host correction, but the correlation analysis and variability observed would remain unchanged. We note that this approach might not be appropriate for multi-facility analysis, especially in the presence of significantly variable seeing.
@@ 3.1 Optical – gamma-ray flux
The $\gamma$-ray flux and r*-band magnitude correlations are shown in Fig. 1, with each panel showing sources of different classifications (i.e. high-, intermediate-, low- synchrotron peak BL Lac types and FSRQs) and each colour within the panel showing different sources. The data from other object classifications are shown as grey points. There is a clear distinction between the BL Lac and FSRQ sources, with some overlap between the FSRQs and BL Lac LSPs. The FSRQ sources generally have higher $\gamma$-ray fluxes; however, this may just be due to the biased nature of the sample selection (sources were added to the observing campaign if they showed high levels of $\gamma$-ray activity with Fermi).
Figure 1: Fermi $\gamma$-ray flux vs. optical r* magnitude for the sample. Each panel highlights the data for the different blazar subclasses separately with the upper left, upper right, lower left, and lower right highlighting data for FSRQs, BL Lac LSPs, ISP BL Lacs, and HSP BL Lacs respectively. In each panel, the data for the other subclasses is shown as faint grey circles.
Using the redshifts, $z$, from Table 1, the $\gamma$-ray fluxes and optical magnitudes can be calibrated for distance. To do this the luminosity distance, $d_{\mathrm{L}}$, was calculated for each object using the WMAP9 cosmology module in Astropy. This module assumes a flat universe, with a Hubble constant of $H_{0}=69.32\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and the matter density parameter set at $\Omega_{\mathrm{m}}=0.2865$ (Hinshaw et al., 2013).
@@
from Hovatta et al. (2014) where $\Gamma$ is the power-law index (taken from the Fermi LCR), F is the $\gamma$-ray flux given in $\mathrm{erg}\,cm^{-2}\,s^{-1}$ in the 1–100 GeV photon energy range, and $d_{\mathrm{L}}$ is the luminosity distance given in cm. The absolute magnitude was calculated by
@@
Figure 2: As Fig. 1, but with Fermi $\gamma$-ray flux and r* magnitude calibrated for distance ($\gamma$-ray luminosity and r* absolute magnitude respectively). Additionally, redshift vectors are included to show the direction and distance the data for a given source would move if the redshift value given in Table 1 was incorrect. The decreasing line indicates we’ve overestimated the redshift by double and the increasing underestimated by half.
We note that by definition, BL Lac objects have difficult-to-determine redshifts due to the comparably small (or absent) emission lines in their optical spectra. This means the calculation of absolute magnitude and $\gamma$-ray luminosity is subject to this uncertainty. To account for this, 1000 redshift values between $\frac{z}{2}<z<2z$ (where $z$ is the redshift value stated in Table 1) were used to calculate the average absolute magnitude and $\gamma$-ray luminosity as functions of redshift. These were plotted and fitted showing the data followed a linear relationship of the form $y=Ax+B$ where $A=-0.504\pm 0.012$ and $B=33.90\pm 0.20$ across all sources. Furthermore, the vector distance the data shifted was the same for redshift values $\frac{z}{2}$ and $2z$ at $1.971\pm 0.023$ units in Fig. 2. Therefore, redshift vectors in the upper left corner of each panel of Fig. 2 are displayed. These vectors show the distance and direction the data for a given source would shift for a 50 per cent reduction ($\frac{z}{2}$) and a 100 per cent increase ($2z$) in the redshift value stated in Table 1. If the true redshift was half the stated value, implying an overestimation, the object would be intrinsically fainter and so would shift towards the point (0,33.90) in Fig. 2. Conversely, if the true redshift was double the stated value, implying an underestimation, the object would be intrinsically brighter and would shift away from the point (0,33.90) in Fig. 2.
Table 5 in the appendix shows the Spearman rank $p$ values and coefficients ($c$) for correlations between the RINGO3 wavebands and $\gamma$-ray flux for each source in the sample; excluding IC310, 1ES 1426+428, and MRK421 for the reasons discussed previously. Of these 28 sources, 21 showed significant correlations between $\gamma$-ray flux and each optical b*g*r* magnitude. All significant correlations were positive. Breaking these correlations down by subclassification we find that 33 per cent of HSP BL Lac sources showed significant correlations, increasing to 66 per cent for ISP BL Lac sources. All BL Lac LSP sources showed significant correlations between the optical and $\gamma$-ray fluxes, along with 85 per cent of FSRQs.
@@ 3.2 Optical spectral index – flux
$F_{\nu}\propto\nu^{-\alpha}$ (3)
where $F_{\nu}$ is the flux at wavelength $\nu$, $\nu$ is the central wavelength of the RINGO3 bands, and $\alpha$ is the spectral index. A two-point spectral index was calculated by taking the gradient of log flux vs. log frequency, giving $-\alpha$. This was done using the r* and b* data, and by fitting a linear least-squares regression. It is important to not include the g*-band data in this calculation to not induce false correlations in the subsequent analysis arising from correlated errors (see Appendix A for more details). The uncertainty of the spectral index at each epoch was calculated using Monte Carlo resampling where, at each epoch, 1000 pairs of randomly generated r* and b* flux values within the respective error limits were generated and the spectral index was calculated. The standard error on these 1000 values was taken as the error on the spectral index.
@@
Table 6 in the appendix shows the results of the above analysis, with the ‘Fit’ column describing which model fits the data better according to the AIC and BIC values. We note that there were no cases with conflicting AIC and BIC values. The ‘Trend’ column describes the colour relationship observed given the obtained preferred fit and sign on the Spearman rank strength coefficient, $c$. Negative strengths indicate bluer-when-brighter (BWB) behaviour, implying the spectral index flattens during periods of heightened flux. Conversely, positive strength coefficients indicate redder-when-brighter (RWB) behaviour implying the spectral index steepens during periods of heightened flux. In both cases, where a log fit is preferred over a linear one, the behaviour becomes more stable as brightness increases meaning the rate at which the colour changes decreases, or altogether flattens; bluer-stable-when-brighter and redder-stable-when-brighter (BSWB and RSWB, respectively). The preferred fit for each source is included in Fig. 3. A dotted line indicates the linear fit and a solid line indicates a logarithmic fit.
@@
Figure 3: Spectral index vs. optical g*-band flux for each object in our sample. A best-fit line is fitted to each set of data after having linear and logarithmic fits qualitatively assessed using AIC and BIC coefficients. A preferential linear fit is shown by a dotted fit line, while a logarithmic fit is shown by a solid line. The Spearman rank correlation coefficients associated with each correlation (linearised in the case of a logarithmic fit) are given in Table 6.